# Beyond the Eye: A Relational Model for Early Dementia Detection Using Retinal OCTA Images

### v1

@@ Abstract
Early detection of dementia, such as Alzheimer’s disease (AD) or mild cognitive impairment (MCI), is essential to enable timely intervention and potential treatment. Accurate detection of AD/MCI is challenging due to the high complexity, cost, and often invasive nature of current diagnostic techniques, which limit their suitability for large-scale population screening. Given the shared embryological origins and physiological characteristics of the retina and brain, retinal imaging is emerging as a potentially rapid and cost-effective alternative for the identification of individuals with or at high risk of AD. In this paper, we present a novel PolarNet+ that uses retinal optical coherence tomography angiography (OCTA) to discriminate early-onset AD (EOAD) and MCI subjects from controls. Our method first maps OCTA images from Cartesian coordinates to polar coordinates, allowing approximate sub-region calculation to implement the clinician-friendly early treatment of diabetic retinopathy study (ETDRS) grid analysis. We then introduce a multi-view module to serialize and analyze the images along three dimensions for comprehensive, clinically useful information extraction. Finally, we abstract the sequence embedding into a graph, transforming the detection task into a general graph classification problem. A regional relationship module is applied after the multi-view module to excavate the relationship between the sub-regions. Such regional relationship analyses validate known eye-brain links and reveal new discriminative patterns. The proposed model is trained, tested, and validated on four retinal OCTA datasets, including 1,671 participants with AD, MCI, and healthy controls. Experimental results demonstrate the performance of our model in detecting AD and MCI with an AUC of 88.69% and 88.02%, respectively. Our results provide evidence that retinal OCTA imaging, coupled with artificial intelligence, may serve as a rapid and non-invasive approach for large-scale screening of AD and MCI.
@@
The eye and brain share similar tissue origin, and their similarity and association of structural characteristics and functional mechanisms have been previously investigated [42, 24]. Recently, significant differences in retinal biomarkers between AD and healthy controls have been reported, suggesting that AD affects the eyes and leads to changes in retinal structures. For example, Wu et al.[38] reported that both AD and MCI patients have a loss of retinal microvascular density in the macular region. Curcio [10] extracted structural features in optical coherence tomography angiography (OCTA), and showed that cognitively impaired participants have a significant decrease in the thickness of inner fovea compared to healthy controls. Zabel et al. [43] demonstrated that AD is associated with retinal neuronal apoptosis and retinal vascular dysfunction. These studies suggest that retinal microvascular attenuation may serve as a potential biomarker for signs of MCI and AD. Thus, retinal imaging has become a potential tool for detecting AD and MCI, and previous studies [20, 7] have mainly applied machine learning methods over color fundus photography (CFP) for AD/MCI detection. Kim et al. [20] used CFP to train a modified MobileNet model to identify individuals with AD. This work modified the attention mechanism to the weighted attention mechanism and applied the mask-adding process to predict the likelihood of AD. Cheung et al. [7] used thousands of CFP samples to develop a deep-learning model for the detection of AD. These methods rarely follow the clinical region-based analysis routine, which limits their ability to incorporate valuable clinical statistical findings and generate easily interpretable results.
@@
Most existing work has used the CFP imaging modality for the study of AD or MCI. Although CFP has advantages in terms of accessibility and cost-effectiveness, its native resolution (60-300 in vessel diameter) is insufficient to image the retinal microvascular network in detail, as demonstrated in Fig. 1-(a): this prevents the detection of subtle vascular changes in the early stages of diseases such as MCI. In contrast, Optical Coherence Tomography Angiography (OCTA) is an advanced imaging modality that provides non-invasive and rapid imaging of the retinal microvasculature and choroidal capillaries with high resolution (5-6 in diameter) across multiple layers, including the superficial vascular complex (SVC), deep vascular complex (DVC) and choriocapillaris (CC). Their maximum projection of OCTA flow signals, a.k.a. en face images (Fig. 1-(b)), which enhanced depth-resolved microvascular imaging facilitates the detection of subtle vascular changes, is critical for accurate screening and early diagnosis of many eye-related diseases, such as diabetic retinopathy [31].
@@
In clinical practice, ophthalmologists often use regional analysis tools to study retinal biomarkers on retinal images. The most commonly used is the Early Treatment of Diabetic Retinopathy Study grid (ETDRS), as shown in Fig. 1-(c). The ETDRS grid is a standardized grid that divides the retina into nine regions with three concentric circles and two orthogonal lines: a central foveal ring of 1 mm diameter, an inner macular ring, and an outer macular ring. The ETDRS grid provides a systematic and consistent assessment of the macular region, allowing a more specific evaluation of retinal changes in a standardized manner, which can provide a more nuanced understanding of the disease[30, 11, 41]. As illustrated in Fig. 1-(c), it can be found that both the ETDRS grid and the en face images of the retina have circular characteristics that follow the nature of biology. Large vessels in SVC grow around a circle and gradually disappear near the center of the circle, forming a capillary-free foveal avascular zone (FAZ)[8]. Thus, in our previous work [40], we used the ETDRS grid to investigate the association of structural features with AD and MCI, and the results have shown significant reductions in vessel area density and vessel length density in specific regions of the inner vascular complexes in AD and MCI participants.
@@
Inspired by the above observations and findings, we propose a novel end-to-end framework in this work, namely PolarNet+, to take full advantage of clinical region-based analysis for EOAD and MCI detection using OCTA images. We aim to integrate the region-based feature extraction procedure, which is consistent with the ETDRS grid, into a deep learning-based classification model. To obtain a more accurate and interpretable result, it is worth noting that we specifically designed an approximate sector convolution, and the polar transformation was then applied to the regions of the ETDRS grid in order to take advantage of spatial constraints and improve the feature extraction and classification performances.
@@
The proposed PolarNet+ significantly extends our work published in MICCAI-2023[23], which was only verified on a single dataset containing 114 subjects. In this work, we expand the data pool from 114 to 1671 participants and make significant technical improvements: as a result, the new approach can also discriminate MCI subjects from healthy controls using retinal OCTA images, while achieving more accurate detection results. Overall, the major differences with the previous work and the main contributions of this work can be summarised as follows:
@@
We consider our work to be the first attempt in this research area for the detection of EOAD and MCI using retinal OCTA images. In addition to providing classification results, our model also provides a regional importance map and a regional relationship graph, highlighting discriminative patterns that drive decision-making, and revealing connections between retinal regions that are informative of neurological disease, respectively.
@@
We introduce a rewiring-based graph reasoning module to fully exploit and understand the relationships across diverse retinal and choroidal layers, leveraging the unique characteristics of OCTA data. This achieves promising classification performances. More importantly, it provides interpretable region connections to improve model transparency.
Our interpretability analyses validate known eye-brain links and reveal new discriminative patterns, demonstrating the model’s potential as a computer-aided pathology tool for studying little-known associations between ophthalmic and complex neurological conditions.
@@ 2 METHOD
Fig. 2 illustrates the outline of our EOAD/MCI detection method using multiple en face angiograms of OCTA as input. We first employ VAFF-Net[16] to locate the FAZ center point on the SVC layer, and then transform the original images from polar coordinates to Cartesian coordinates, as shown in Fig. 2 (a) and (b). These transformed images are fed into the PolarNet+ for the extraction of sequential features in circle-area, ring-area, and sector-area along three dimensions. After sequencing, the sequences that encode complementary region-specific information are treated as graph nodes and fed into a regional relationship module. Finally, PolarNet+ aggregates the node features and relationships for the generation of the final detection output. In addition, PolarNet+ can provide two visualizations for improved model interpretability, as demonstrated in Fig. 2 (e) and (f): a region importance map highlighting discriminative patterns that drive the decision-making, and a regional relationship graph revealing connections between retinal areas that are informative for neurological conditions.
@@ 2.1 Polar transformation of OCTA image
Here we introduce the polar transformation to realize region-based analysis. It aims to map coordinates from the polar coordinates to the Cartesian coordinate system . Many image analysis tasks, such as pattern analysis, shape identification, or feature extraction, benefit from this transformation: e.g., Fu et al. [13] use polar transformation to transfer the redial relationship to a spatial relationship on the image segmentation task.
@@ 2.1.1 Approximate sector-shaped computing
For the sake of simplicity and clarity, Fig. 2(c) illustrates the ETDRS grids mapped on the polar transformed image. Approximate sector calculation functions can be adapted to any sector herein. For instance, when we perform a rectangular kernel convolution along the direction on the transformed image, it is equivalent to performing a sector kernel convolution around the center of the FAZ in the original image in a counterclockwise direction.
@@ 2.2 The classification model: PolarNet+
OCTA provides in-depth information to visualize the retinal and choroidal microvascular network and the structure of the FAZ at different layers. To this end, we take advantage of three different en face angiograms from different layers and explore intra- and inter-instance relationships across different retinal and choroidal layers, exploiting the unique characteristics of OCTA data. In this work, the SVC, DVC, and CC layers are used as input to the given PolarNet+, and we stack these three en face images as 3D volume data.
@@ 2.2.1 Spatial Extension Module (SEM)
Denote the input OCTA image as of size , , and K as the total number of images from one patient. The spatial extension module consists of a multi-layer perceptrons (MLP) layer, a convolution layer, and a ReLU activation layer, which can be represented as:
@@ 2.2.3 Regional Relationship Module (RRM)
Modeling the relationships and connections between different regions in the ETDRS grid is critical to establishing the association between pathological conditions and the retinal microvasculature for the accurate detection of EOAD and MCI. To this end, we propose the regional relationship module (RRM), which transforms the image classification problem into a graph representation, with region sequences modeled as graph nodes. Graph neural networks are naturally suited to learning from relational data and can embed high-dimensional connectivity patterns in a low-dimensional space [39]. Critically, rather than relying on a predefined adjacency matrix, the RRM incorporates a rewiring mechanism[3] to learn the edges between regions. Through rewiring, it can learn an enriched representation of the relationships [28] between different regions within the ETDRS grid, which then allows the revelation of critical interdependencies associated with disease. The RRM contains two differentiable layers, the commute times layer and the spectral gap optimization layer, for rewiring. The commute times layer [3] identifies salient edges based on resistance; while the spectral gap optimization [3] layer optimizes the topology using spectral graph theory.
To find optimal clusters in the graph, we introduce MinCUT pooling [6]. Given an adjacency matrix and a feature matrix , we get a node cluster where denotes the number of nodes and denotes the dimensions of the features. For a given input node cluster , a MinCUT pooling layer is applied to learn a new number of clusters: .
@@ 3 Datasets
Two different datasets were used: ROAD (Retinal OCTA for EOAD study) and ROMCI (Retinal OCTA for MCI study). All OCTA images in these two datasets were obtained from a multi-center case-control study for the detection of AD and MCI. The SVC, DVC, and CC angiograms were used to train, test, and validate the proposed automated method, because together they contain the complete vasculature from superficial to deep. Images from any given subject were alloted only to either the training or test sets, to avoid information leakage. All images analyzed were fovea-centered. The clinical protocol of this study was approved by the Ethics Committee of the Cixi Institute of Biomedical Engineering, Chinese Academy of Sciences, and adhered to the tenets of the Declaration of Helsinki. Informed written consent was obtained from the participants enrolled in our study.
3.0.1 Retinal OCTA data for EOAD study (ROAD)
@@
3.0.2 Retinal OCTA data for MCI study (ROMCI)
@@ 4.1 Implementation details
We selected SVC, DVC, and CC as the input images, and the polar transformation was applied to them. Because the ETDRS regions of two eyes are symmetrical, before the transformation, all the images were flipped from the left eye (OS) to the right eye (OD). The transformed images were scaled down to a width of 224 pixels. We also applied data augmentation by randomly rotating the images by 20 degrees around the transformation centers (Fig. 2(a)). We implemented our method with PyTorch. A server running Ubuntu 20.04 with two Nvidia RTX 3090 GPUs was used to train the model. An initial learning rate of 2e-5 and a batch size of 16 were used, with AdamW [27] as the optimizer. Fivefold cross-validation was used to exploit the data and maximise reliability. It is worth noting that the proposed PolarNet+ was implemented by repeating the Res3D block and the Sequencer3D blocks twice each to avoid over-fitting during training.
@@ 4.3.1 Regional importance analysis
In the testing stage, we engaged the PRIM and produced a importance matrix, and accumulated the entire testing set. Subsequently, we performed an inverse operation of the polar transformation to generate the corresponding importance map based on the ETDRS grid, as shown in Fig. 4-(a). Furthermore, we analyzed eight distinct parameters that collectively described the retinal microvasculature and the FAZ. These parameters encompass vascular length density (VLD), vascular area density (VAD), vascular bifurcation number (VB), vascular fractal dimension (VFD), FAZ area (FA), FAZ circularity (FC), FAZ roundness (FR), and FAZ solidity (FS). Subsequently, we examined the disparities between the EOAD/MCI and control groups, with the findings presented in Fig. 4-(c). Compared to a commonly-used visualization method, such as Grad-CAM [33], as shown in Fig. 4-(b), our region importance map is more clinician-friendly and more easily understandable.

### v2

@@ Abstract
Early detection of dementia, such as Alzheimer’s disease (AD) or mild cognitive impairment (MCI), is essential to enable timely intervention and potential treatment. Accurate detection of AD/MCI is challenging due to the high complexity, cost, and often invasive nature of current diagnostic techniques, which limit their suitability for large-scale population screening. Given the shared embryological origins and physiological characteristics of the retina and brain, retinal imaging is emerging as a potentially rapid and cost-effective alternative for the identification of individuals with or at high risk of AD. In this paper, we present a novel PolarNet+ that uses retinal optical coherence tomography angiography (OCTA) to discriminate early-onset AD (EOAD) and MCI subjects from controls. Our method first maps OCTA images from Cartesian coordinates to polar coordinates, allowing approximate sub-region calculation to implement the clinician-friendly early treatment of diabetic retinopathy study (ETDRS) grid analysis. We then introduce a multi-view module to serialize and analyze the images along three dimensions for comprehensive, clinically useful information extraction. Finally, we abstract the sequence embedding into a graph, transforming the detection task into a general graph classification problem. A regional relationship module is applied after the multi-view module to explore the relationship between the sub-regions. Such regional relationship analyses validate known eye-brain links and reveal new discriminative patterns. The proposed model is trained, tested, and validated on four retinal OCTA datasets, including 1,671 participants with AD, MCI, and healthy controls. Experimental results demonstrate the performance of our model in detecting AD and MCI with an AUC of 88.69% and 88.02%, respectively. Our results provide evidence that retinal OCTA imaging, coupled with artificial intelligence, may serve as a rapid and non-invasive approach for large-scale screening of AD and MCI. The code is available at https://github.com/iMED-Lab/PolarNet-Plus-PyTorch, and the dataset is also available upon request.
@@
The eye and brain share a similar tissue origin, and their similarity and association of structural characteristics and functional mechanisms have been previously investigated [44, 25]. Recently, significant differences in retinal biomarkers between AD and healthy controls have been reported, suggesting that AD affects the eyes and leads to changes in retinal structures. For example, Wu et al.[39] reported that both AD and MCI patients have a loss of retinal microvascular density in the macular region. Curcio [10] extracted structural features in optical coherence tomography angiography (OCTA), and showed that cognitively impaired participants have a significant decrease in the thickness of inner fovea compared to healthy controls. Zabel et al. [45] demonstrated that AD is associated with retinal neuronal apoptosis and retinal vascular dysfunction. These studies suggest that retinal microvascular attenuation may serve as a potential biomarker for signs of MCI and AD. Thus, retinal imaging has become a potential tool for detecting AD and MCI, and previous studies [21, 7] have mainly applied machine learning methods over color fundus photography (CFP) for AD/MCI detection. Kim et al. [21] used CFP to train a modified MobileNet model to identify individuals with AD. This work modified the attention mechanism to the weighted attention mechanism and applied the mask-adding process to predict the likelihood of AD. Cheung et al. [7] used thousands of CFP samples to develop a deep-learning model for the detection of AD. These methods rarely follow the clinical region-based analysis routine, which limits their ability to incorporate valuable clinical statistical findings and generate clinical-friendly results.
@@
Most existing work has used the CFP imaging modality for the study of AD or MCI. Although CFP has advantages in terms of accessibility and cost-effectiveness, its native resolution (60-300 in vessel diameter) is insufficient for imaging the retinal microvascular network in detail, as demonstrated in Fig. 1-(a): this prevents the detection of subtle vascular changes in the early stages of diseases such as MCI. In contrast, Optical Coherence Tomography Angiography (OCTA) is an advanced imaging modality that provides non-invasive and rapid imaging of the retinal microvasculature, and choroidal capillaries with high resolution (5-6 in diameter) across multiple layers, including the superficial vascular complex (SVC), deep vascular complex (DVC) and choriocapillaris (CC). Their maximum projection of OCTA flow signals, a.k.a. en face images (Fig. 1-(b)), which enhances depth-resolved microvascular imaging and facilitates the detection of subtle vascular changes, is critical for accurate screening and early diagnosis of many eye-related diseases, such as diabetic retinopathy [32].
@@
In clinical practice, ophthalmologists often use regional analysis tools to study retinal biomarkers on retinal images. The most commonly used is the Early Treatment of Diabetic Retinopathy Study grid (ETDRS), as shown in Fig. 1-(c). The ETDRS grid is a standardized grid that divides the retina into nine regions with three concentric circles and two orthogonal lines: a central foveal ring of 1 mm diameter, an inner macular ring, and an outer macular ring. The ETDRS grid provides a systematic and consistent assessment of the macular region, allowing a more specific evaluation of retinal changes in a standardized manner, which can provide a more nuanced understanding of the disease[31, 11, 42]. As illustrated in Fig. 1-(c), it can be found that both the ETDRS grid and the en face images of the retina have circular characteristics that follow the nature of biology. Large vessels in SVC grow around a circle and gradually disappear near the center of the circle, forming a capillary-free foveal avascular zone (FAZ)[8]. Thus, in our previous work [41], we used the ETDRS grid to investigate the association of structural features with AD and MCI, and the results showed significant reductions in vessel area density and vessel length density in specific regions of the inner vascular complexes in AD and MCI participants.
@@
Inspired by the above observations and findings, we propose a novel end-to-end framework in this work, namely PolarNet+, to take full advantage of clinical region-based analysis for EOAD and MCI detection using OCTA images. We aim to integrate the region-based feature extraction procedure, which is consistent with the ETDRS grid, into a deep learning-based classification model. To obtain a more accurate and clinical-acceptable result, it is worth noting that we specifically designed an approximate sector convolution, and the polar transformation was then applied to the regions of the ETDRS grid in order to take advantage of spatial constraints and improve the feature extraction and classification performances.
@@
The proposed PolarNet+ significantly extends our work published in MICCAI-2023[24], which was only verified on a single dataset containing 114 subjects. In this work, we expand the data pool from 114 to 1671 participants and make significant technical improvements. As a result, the new approach can also discriminate MCI subjects from healthy controls using retinal OCTA images, while achieving more accurate detection results. Overall, the major differences with the previous work and the main contributions of this work can be summarised as follows:
@@
We consider our work to be the first attempt in this research area for the detection of EOAD and MCI using retinal OCTA images. In addition to providing classification results, our model also provides a regional importance map and a regional relationship graph, highlighting discriminative patterns that drive decision-making, and revealing connections between retinal regions that are informative of neurological disease.
@@
We introduce a rewiring-based graph reasoning module to fully exploit and understand the relationships across diverse retinal and choroidal layers, leveraging the unique characteristics of OCTA data. This achieves promising classification performance. More importantly, it provides region connections to improve model transparency thereby facilitating a more clinically acceptable insight.
Our clinical-friendly analyses validate known eye-brain links and reveal new discriminative patterns, demonstrating the model’s potential as a computer-aided pathology tool for studying little-known associations between ophthalmic and complex neurological conditions.
@@ 2 METHOD
Fig. 2 illustrates the outline of our EOAD/MCI detection method using multiple en face angiograms of OCTA as input. We first employ VAFF-Net[17] to locate the FAZ center point on the SVC layer, and then transform the original images from polar coordinates to Cartesian coordinates, as shown in Fig. 2 (a) and (b). These transformed images are fed into the PolarNet+ for the extraction of sequential features in circle-area, ring-area, and sector-area along three dimensions. After sequencing, the sequences that encode complementary region-specific information are treated as graph nodes and fed into a regional relationship module. Finally, PolarNet+ aggregates the node features and relationships for the generation of the final detection output. Furthermore, PolarNet+ is capable of generating two distinct visualisations, thereby facilitating a more clinically acceptable insight., as demonstrated in Fig. 2 (e) and (f): a region importance map highlighting discriminative patterns that drive the decision-making, and a regional relationship graph revealing connections between retinal areas that are informative for neurological conditions.
@@ 2.1 Polar transformation of OCTA image
Here we introduce the polar transformation to realize region-based analysis. It aims to map coordinates from the polar coordinates to the Cartesian coordinate system . Many image analysis tasks, such as pattern analysis, shape identification, or feature extraction, benefit from this transformation: e.g., Fu et al. [13] use polar transformation to convert the radial relationship to a spatial relationship in the image segmentation task.
@@ 2.1.1 Approximate sector-shaped computing
For the sake of simplicity and clarity, Fig. 2(c) illustrates the ETDRS grids mapped on the polar transformed image. Approximate sector calculation functions can be adapted to any sector herein. For instance, when we perform a rectangular kernel convolution along the direction on the transformed image, it is equivalent to performing a sector-shaped kernel convolution around the center of the FAZ in the original image in a counterclockwise direction.
@@ 2.2 The classification model: PolarNet+
OCTA provides in-depth information to visualize the retinal and choroidal microvascular network and the structure of the FAZ at different layers. To this end, we take advantage of three different en face angiograms from different layers and explore intra-instance and inter-instance relationships across different retinal and choroidal layers, exploiting the unique characteristics of OCTA data. In this work, the SVC, DVC, and CC layers are used as input to the given PolarNet+, and we stack these three en face images as 3D volume data.
@@ 2.2.1 Spatial Extension Module (SEM)
Denote the input OCTA image as of size , , and K as the total number of images from one patient. The spatial extension module consists of a multi-layer perceptron (MLP) layer, a convolution layer, and a ReLU activation layer, which can be represented as:
@@ 2.2.3 Regional Relationship Module (RRM)
Modeling the relationships and connections between different regions in the ETDRS grid is critical to establishing the association between pathological conditions and the retinal microvasculature for the accurate detection of EOAD and MCI. To this end, we propose the regional relationship module (RRM), which transforms the image classification problem into a graph representation, with region sequences modeled as graph nodes. Graph neural networks are naturally suited to learning from relational data and can embed high-dimensional connectivity patterns in a low-dimensional space [40]. Critically, rather than relying on a predefined adjacency matrix, the RRM incorporates a rewiring mechanism [3] to learn the edges between regions. Through rewiring, it can learn an enriched representation of the relationships [29] between different regions within the ETDRS grid, which then allows the revelation of critical interdependencies associated with disease. The RRM contains two differentiable layers, the commute times layer and the spectral gap optimization layer, for rewiring. The commute times layer [3] identifies salient edges based on resistance, while the spectral gap optimization [3] layer optimizes the topology using spectral graph theory.
To find optimal clusters in the graph, we introduce MinCUT pooling [6]. Given an adjacency matrix and a feature matrix , we obtain a node cluster where denotes the number of nodes and denotes the dimensions of the features. For a given input node cluster , a MinCUT pooling layer is applied to learn a new number of clusters: .
@@ 3 Datasets
Two different datasets were used: ROAD (Retinal OCTA for EOAD study) and ROMCI (Retinal OCTA for MCI study). All OCTA images in these two datasets were obtained from a multi-center case-control study for the detection of AD and MCI. The SVC, DVC, and CC angiograms were used for training, validation, and testing of the proposed automated method, as they together provide a comprehensive representation of the vasculature from superficial to deep. Images from any given subject were allocated only to either the training or test sets, to avoid information leakage. All images analyzed were fovea-centered. The clinical protocol of this study was approved by the Ethics Committee of the Cixi Institute of Biomedical Engineering, Chinese Academy of Sciences, and adhered to the tenets of the Declaration of Helsinki. Informed written consent was obtained from the participants enrolled in our study.
3.1 Retinal OCTA data for EOAD study (ROAD)
@@
3.2 Retinal OCTA data for MCI study (ROMCI)
@@ 4.1 Implementation details
To mitigate the impact of data imbalance, we employed the following strategies. Firstly, we applied data augmentation techniques, including rotating the minority class images (±20 degrees), to increase the diversity and representation of EOAD and MCI samples. In addition, a class-weighted loss function was used during model training to ensure balanced attention to all classes, regardless of their frequency in the dataset. The same data augmentation procedures, including rotation by +/- 20 degrees, were applied consistently across all comparison methods to ensure a fair comparison.
We selected SVC, DVC, and CC as the input images, and the polar transformation was applied to them. Because the ETDRS regions of two eyes are symmetrical, before the transformation, all the images were flipped from the left eye (OS) to the right eye (OD). The transformed images were scaled down to a width of 224 pixels. We performed a hyperparameter search for all methods involved in the comparison (initial learning rate: from 1e-5 to 2e-4, stepped by 3e-5; dropout rate: from 4e-2 to 2e-1, stepped by 2e-2), and applied the corresponding hyperparameters that yielded optimal performance. The batch size is fixed to 16 and the same data augmentation procedures were applied consistently across all methods to ensure a fair comparison.
We implemented our method with PyTorch. A server running Ubuntu 20.04 with two Nvidia RTX 3090 GPUs was used to train the model. An initial learning rate of 2e-5 and a batch size of 16 were used, with AdamW [28] as the optimizer. Fivefold cross-validation was used to exploit the data and maximise reliability. It is worth noting that the proposed PolarNet+ was implemented by repeating the Res3D block and the Sequencer3D blocks twice each to avoid over-fitting during training.
@@ 4.3.1 Regional importance analysis
In the testing stage, we engaged the PRIM and produced a importance matrix, and accumulated the entire testing set. Subsequently, we performed an inverse operation of the polar transformation to generate the corresponding importance map based on the ETDRS grid, as shown in Fig. 5-(a). Furthermore, we analyzed eight distinct parameters that collectively described the retinal microvasculature and the FAZ. These parameters encompass vascular length density (VLD), vascular area density (VAD), vascular bifurcation number (VB), vascular fractal dimension (VFD), FAZ area (FA), FAZ circularity (FC), FAZ roundness (FR), and FAZ solidity (FS). Subsequently, we examined the disparities between the EOAD/MCI and control groups, with the findings presented in Fig. 5-(c). The statistical analysis has shown that the parameters of VLD, VAD, and VB showed significance in SVC-IE and DVC-II, and VAD and FR showed significance in SVC-SI (all p-values are less than 0.001. Compared to a commonly used visualization method, such as Grad-CAM [34], as shown in Fig. 5-(b), our region importance map is more clinician-friendly and easier to understand.